Find the minimum initial speed of a projectile in order for it to reach a height of 2191 km above the surface of the earth.

________????
Answer #1

the requird height to reach H = 1781km = 1781000m

the required minimum speed v = ?

the expression for height H = v2/2g

v2 = 2gH = 2*9.8*2191000

v = 6553.14 m/s = 6.55 km/s